# A new hierarchy for complex plane curves

### v2

@@ 2.2. Some key exact sequences
Let us come back to the setting of Theorem 1.4. To simplify the notation in the sequel we set
@@
If we identified $D_{0}(f)$ with the quotient $S$-module
@@
where $(f)$ is the principal ideal generated by $f$, and $SE$ denotes the submodule in $D(f)$ generated by the Euler derivation $E=(x,y,z)$. The first nonzero morphism in (2.5) is multiplication by $f_{2}$ if we replace $D_{0}(f_{1})$ and $D_{0}(f)$ by $\overline{D}(f_{1})$ and $\overline{D}(f)$ respectively, see for instance [28, Theorem 1.6]. This fact is used in the proof of Theorems 1.19 and 1.24 below.
@@ 2.4. Exact sequences for line arrangements
Consider the triple of line arrangements $({\mathcal{A}}^{\prime},{\mathcal{A}},{\mathcal{A}}^{\prime\prime})$, where ${\mathcal{A}}^{\prime}={\mathcal{A}}\setminus\{L\}$, $L$ being a line in ${\mathcal{A}}$, and ${\mathcal{A}}^{\prime\prime}$ is the arrangement consisting of points on the line $L$ corresponding to the intersection set ${\mathcal{A}}^{\prime}\cap L$. By a change of coordinates we can take $L:x=0$ and set
${\mathcal{A}}^{\prime}:f^{\prime}=0,\ {\mathcal{A}}:f=0\text{ and }{\mathcal{A}}^{\prime\prime}:f^{\prime\prime}=0,$
@@
where $D(f^{\prime})$, $D(f)$ and $D(f^{\prime\prime})$ are as in Remark 2.3 and in addition $D(f^{\prime\prime})$ is regarded as an $S$-module using the restriction of scalars given by the ring morphism $\psi:S\to R$, which replaces $x$ by 00 in any polynomial $h\in S$. Here the first morphism $D(f^{\prime})(-1)\to D(f)$ is the multiplication by $x$, and the second morphism is given by
@@
see [25, Proposition 4.45] and [27, Lemma 2.6] that covers a more general case when ${\mathcal{A}}^{\prime}$ is a conic-line arrangement, maybe with some quasi-homogeneity conditions. If we replace the modules $D(f^{\prime})$, $D(f)$ and $D(f^{\prime\prime})$ by the modules $\overline{D}(f^{\prime})$, $\overline{D}(f)$ and $\overline{D}(f^{\prime\prime})$, the sequence obtained from (2.8) is no longer exact. However, if we pass to the associated coherent sheaves on $\mathbb{P}^{2}$, we get the following short exact sequence
@@
Let $|{\mathcal{A}}^{\prime\prime}|=n$, and note that this implies $|{\mathcal{A}}^{\prime}|\geq n$. Then the $S$-module $D({\mathcal{A}}^{\prime\prime})$ has two generators, one of degree $1$ and another of degree $n-1$. The first element is the Euler derivation, which is in the image of $\rho$ always. It follows that
@@
is surjective. Moreover, it suffices to show that $\overline{\rho}$ is surjective in degree $n-1$. Using the exact sequence (2.9), it suffices to show that $H^{1}(\mathbb{P}^{2},E_{{\mathcal{A}}^{\prime}}(n-2))=0$. By the minimal free resolution of $D_{0}({\mathcal{A}}^{\prime})$, we have
00 0 $=$ $H^{1}(\mathbb{P}^{2},\oplus\mathcal{O}(-d_{i}+n-2))\rightarrow H^{1}(\mathbb{P}^{2},E_{{\mathcal{A}}^{\prime}}(n-2))$
@@
The exact sequence in Theorem 2.5 plays a key role in our proof of Theorem 1.24. When we pass from line arrangements to conic-line arrangements, there are similar exact sequences according to [27], and they are discussed and used in the proof of Theorem 1.19 below.
3. The proofs of Theorems 1.4 and 1.7 and of Corollaries 1.9 and 1.11
3.1. The proof of Theorem 1.4
There is a natural injective $S$-linear morphism
@@
where $d=\deg C=e_{1}+e_{2}$ and $E$ denotes the Euler derivation, see [14, Theorem 5.1]. We denote by $d_{1},d_{2}$ (resp. by $d_{1}^{\prime},d_{2}^{\prime}$ ) the first two exponents of $C_{1}$ (resp. of $C$). By Proposition 1.3, we have
@@
This is a contradiction since the quotient $S$-module $D_{0}(f_{1})/Sr_{1}$ is torsion free, as it follows, for instance, from the exact sequence in [18, Theorem 5.1].
3.2. The proof of Theorem 1.7 and of Corollaries 1.9 and 1.11
@@
and $r=e_{1}e_{2}$. Moreover $D^{\prime}$ consists of $e_{2}$ points, and hence
@@
Now we prove Corollary 1.9. The first claim follows from (1.3), since
@@
Moreover, (1.3) tells us that we can apply Theorem 1.7 to any line arrangement $C_{1}$. Starting with the integers $t$ and $d$ as in the second claim of Corollary 1.9, we choose $C_{1}$ to be a free line arrangement of $e_{1}=d-t\geq 3$ lines, for instance $x^{e_{1}}+y^{e_{1}}=0$. Then we add $t$ generic lines to this arrangement and we get by Theorem 1.7 a line arrangement ${\mathcal{A}}$ of degree $d$ and type $t$. This completes the proof of Corollary 1.9.
To prove Corollary 1.11, we first notice that our additional assumption on ${\mathcal{C}}$ implies
@@
for $e_{1}\geq 4$. To complete the proof of Corollary 1.11 it remains to add $t$ generic lines to the arrangement ${\mathcal{C}}_{1}$ and get in this way the required arrangement ${\mathcal{C}}$ by Theorem 1.7.
@@
The curve ${\mathcal{C}}$ has a complicated singularity at the point $p=(0:0:1)$. Indeed, looking at $({\mathcal{C}},p)$ as the union of a $(e_{1}-2)$-ordinary point with the smooth germ at $p$ given by the conic $xz+y^{2}=0$ and using the known formula for the Milnor number of a union, see for instance [30, Theorem 6.5.1], one can show that the Milnor number $\mu({\mathcal{C}},p)$ is given by
@@
which was introduced in [8] and which measure how far is $({\mathcal{C}},p)$ from being quasi-homogeneous, is given by
@@
4. The proof of Propositions 1.13, 1.14 and 1.15
4.1. The proof of Proposition 1.13
To see that the number $m$ of syzygies can be only $3$ or $4$, and to get the corresponding values for the $\varepsilon_{j}$, use the equality (1.7). In case (1), the value for $\tau(C)$ follows from [17, Proposition 2.1] using that $d_{1}+d_{2}=d+1$, where $d=\deg C$. In case (2), we use (1.6) to get the following resolution of the Milnor algebra $M(f)$:
@@
4.2. The proof of Propositions 1.14 and 1.15
We start with Proposition 1.14, and use Theorem 2.1 and Proposition 1.13. Note that the condition $d_{1}<d_{2}-2$ is exactly the condition $d_{1}<(d-1)/2$ which occurs in Theorem 2.1 (1). Hence, to prove the claim $(1)$ in Proposition 1.14, we use the formula for $\nu(C)$ given in Theorem 2.1 (1) and the formula for $\tau(C)$ given in Proposition 1.13, $(1)$. To prove the claims (2), (3) and (4), we use the formula for $\nu(C)$ given in Theorem 2.1 (2) and the formula for $\tau(C)$ given in Proposition 1.13 $(1)$. This completes the proof of Proposition 1.14.
@@
5. The proof of Theorem 1.18, Theorem 1.19, Proposition 1.22 and Theorem 1.24
5.1. The proof of Theorem 1.18
We use the exact sequence (2.6) above with $C_{1}={\mathcal{A}}^{\prime}:f_{1}=0$ and $C_{2}=L:f_{2}=0$. Hence $e_{1}=m_{1}+m_{2}=r$, $e_{2}=1$, $g_{2}=\operatorname{genus}(C_{2})=0$. Moreover $D^{\prime}$ is just a point, and hence
@@
Let $(d_{1},d_{2},...,d_{m})$ be the exponents of the curve $C={\mathcal{A}}$.
Consider first the case $k<m_{1}+m_{2}-1=e_{1}-1$. Then $H^{0}(L,{\mathcal{O}}_{L}(k+1-m_{1}-m_{2}))=0$ and we get an isomorphism
@@
5.2. The proof of Theorem 1.19
This rather long proof is divided into two steps, to make its understanding easier for the reader.
Step 1. We use again the exact sequence (2.6) above with $C_{1}={\mathcal{C}}^{\prime}:f_{1}=0$ and $C_{2}=Q:f_{2}=0$. Hence now we have $e_{1}=m_{1}+m_{2}+1$, $r=2e_{1}$, $e_{2}=2$, $g_{2}=\operatorname{genus}(C_{2})=0$. Moreover, $D^{\prime}$ consists of $2$ points and hence
@@
Let $(d_{1},d_{2},...,d_{m})$ be the exponents of the curve $C={\mathcal{C}}$.
Consider first the case $k<e_{1}$. Then $H^{0}(L,{\mathcal{O}}_{L}(2k-2e_{1}))=0$ and we get an isomorphism
@@
Step 2. If the conic-line arrangement ${\mathcal{C}}$ has type $2A$, then the exponents of ${\mathcal{C}}$ are $(d_{1},d_{2},d_{3})=(m_{1}+2,m_{2}+2,m_{1}+m_{2}+1)$, and our result is proved. To show that this is the case, we have to show that there is no need of an additional generator for $D_{0}(f)$ beyond ${\alpha}$.
To do this, we recall some key results from [27]. A triple $({\mathcal{C}}^{\prime},{\mathcal{C}},{\mathcal{C}}^{\prime\prime})$ of CL (short for conic-line) arrangements, where ${\mathcal{C}}^{\prime}={\mathcal{C}}\setminus\{Q\}$ with $Q\in{\mathcal{C}}$ a smooth conic and ${\mathcal{C}}^{\prime\prime}$ the 0-dimensional arrangement on $Q$ given by the reduced intersection ${\mathcal{C}}^{\prime}\cap Q$, is called quasi-homogeneous if all the singular points of ${\mathcal{C}}^{\prime}$ and of ${\mathcal{C}}$ situated on $Q$ are quasi-homogeneous. The authors in [27] require that all the singular points of ${\mathcal{C}}^{\prime}$ and of ${\mathcal{C}}$ are quasi-homogeneous, but this stronger form is not necessary. Indeed, as noted in [14] and in subsection 2.2 above, for a similar condition in [28], this assumption is used in [27] only to compute the difference
@@
${\mathcal{C}}^{\prime}:f^{\prime}=0\text{ and }{\mathcal{C}}:f=0,$
see for instance the proof of [27, Proposition 2.8 and Proposition3.7]. To do this computation, our weaker assumption is sufficient.
@@
Using this isomorphism, we can regard the 0-dimensional arrangement ${\mathcal{C}}^{\prime\prime}$ as a 0-dimensional arrangement in $\mathbb{P}^{1}$ and hence it is defined by an equation $f^{\prime\prime}=0$, where $f^{\prime\prime}\in R=\mathbb{C}[s,t]$ is some homogeneous polynomial. Hence the graded $R$-module $D(f^{\prime\prime})$ is well defined. Moreover, the map $v$ above induces a ring morphism
@@
and the $S$-module $D(f^{\prime\prime})_{\psi}$ is just the module obtained from the $R$-module $D(f^{\prime\prime})$ by restriction of scalars using $\psi$. Next we consider the morphisms in the exact sequence (5.3). The first morphism $D(f^{\prime})(-2)\to D(f)$ is given by the multiplication by $q$, the equation of $Q$, and hence it preserves the homogeneous components, that is
@@
After these preliminaries, we can continue the proof in Step 2. In our situation, ${\mathcal{C}}^{\prime\prime}$ consists of $2e_{1}$ points, hence $f^{\prime\prime}$ is a homogeneous polynomial of degree $2e_{1}$ without multiple factors. The derivation ${\alpha}\in D_{0}(f)_{e_{1}}$ that was constructed at the end of Step 1 can be regarded as sitting in $\overline{D}(f)_{e_{1}}$ using the identification from Remark 2.3. Then any representative ${\alpha}_{0}\in D(f)_{e_{1}}$ of ${\alpha}$ is clearly not in the image of the morphism $D(f^{\prime})(-2)\to D(f)$ and hence $\rho({\alpha}_{0})\neq 0$. If we assume ${\alpha}_{0}=(a_{1},a_{2},a_{3})$, then
@@
is a new representative of ${\alpha}$ and clearly $\rho({\alpha}_{1})=0$, a contradiction.
If ${\mathcal{C}}$ has type $2B$, it means that there is an additional generator of degree $d_{4}>d_{3}=e_{1}$. Using (1.4), we get $d_{4}\leq(e_{1}+2)-1$, and hence $d_{4}=e_{1}+1$. On the other hand, the exact sequence (5.2) where $e_{1}$ is replaced by $e_{1}+1$ shows that $\overline{D}_{0}(f)_{e_{1}+1}$ modulo the image of $\overline{D}_{0}(f_{1})_{e_{1}-1}$ is 2 dimensional. It follows that we need a new generator of degree $e_{1}+1$ if and only the vector space $V$ spanned by $x{\alpha}_{0},y{\alpha}_{0}$ and $z{\alpha}_{0}$ in $\overline{D}_{0}(f)_{e_{1}+1}$ modulo the image of $\overline{D}(f_{1})_{e_{1}-1}$ is 1-dimensional. Note that $\rho(E)=E^{\prime\prime}$, the Euler derivation corresponding to the ring $R$, and hence there is an induced morphism
@@
5.3. The proof of Proposition 1.22
@@
5.4. The proof of Theorem 1.24
This proof is similar to the proof of Theorem 1.19, so we divide it into two steps.
@@
Let $(d_{1},d_{2},...,d_{m})$ be the exponents of the curve $C={\mathcal{A}}$.

### v3

@@ 2.2. Some key exact sequences
Let us come back to the setting of Theorem 1.4. To simplify the notation, in the sequel we set
@@
If we identify $D_{0}(f)$ with the quotient $S$-module
@@
where $(f)$ is the principal ideal generated by $f$, and $SE$ denotes the submodule in $D(f)$ generated by the Euler derivation $E=(x,y,z)$. The first non-zero morphism in (2.5) is multiplication by $f_{2}$ if we replace $D_{0}(f_{1})$ and $D_{0}(f)$ by $\overline{D}(f_{1})$ and $\overline{D}(f)$ respectively, see for instance [28, Theorem 1.6]. This fact is used in the proof of Theorems 1.16 and 1.20 below.
@@ 2.4. Exact sequences for line arrangements
Consider the triple of line arrangements $({\mathcal{A}}^{\prime},{\mathcal{A}},{\mathcal{A}}^{\prime\prime})$, where ${\mathcal{A}}^{\prime}={\mathcal{A}}\setminus\{L\}$ with $L$ being a line in $\mathcal{A}$}, and $\mathcal{A}$$^{\prime\prime}$ is the arrangement consisting of points on the line $L$ corresponding to the intersection set ${\mathcal{A}}^{\prime}\cap L$. By a change of coordinates we can take $L:x=0$, and we set
[…]
@@
where $D(f^{\prime})$, $D(f)$ and $D(f^{\prime\prime})$ are as in Remark 2.3 and in addition $D(f^{\prime\prime})$ is regarded as an $S$-module using the restriction of scalars given by the ring morphism $\psi:S\to R$, which replaces $x$ by 0 in any polynomial $h\in S$. Here the first morphism $D(f^{\prime})(-1)\to D(f)$ is the multiplication by $x$, and the second morphism is given by
@@
see [25, Proposition 4.45] and [27, Lemma 2.6] that covers a more general case when $\mathcal{A}$$^{\prime}$ is a conic-line arrangement, maybe under some quasi-homogeneity conditions. If we replace the modules $D(f^{\prime})$, $D(f)$ and $D(f^{\prime\prime})$ by the modules $\overline{D}(f^{\prime})$, $\overline{D}(f)$ and $\overline{D}(f^{\prime\prime})$, the sequence obtained from (2.8) is no longer exact. However, if we pass to the associated coherent sheaves on $\mathbb{P}^{2}$, we get the following short exact sequence
@@
Let $|{\mathcal{A}}^{\prime\prime}|=n$, and note that this implies $|{\mathcal{A}}^{\prime}|\geq n$. Then the $S$-module $D({\mathcal{A}}^{\prime\prime})$ has two generators, one of degree $1$ and another of degree n1n-1. The first element is the Euler derivation, which is always in the image of $\rho$. It follows that
@@
is surjective. Moreover, it suffices to show that $\overline{\rho}$ is surjective in degree n1n-1. Using the exact sequence (2.9), it suffices to show that $H^{1}(\mathbb{P}^{2},E_{{\mathcal{A}}^{\prime}}(n-2))=0$. By the minimal free resolution of $D_{0}({\mathcal{A}}^{\prime})$, we have
$0$ $=$ $H^{1}(\mathbb{P}^{2},\oplus\mathcal{O}(-d_{i}+n-2))\rightarrow H^{1}(\mathbb{P}^{2},E_{{\mathcal{A}}^{\prime}}(n-2))$
@@
The exact sequence in Theorem 2.5 plays a key role in our proof of Theorem 1.20. When we pass from line arrangements to conic-line arrangements, there are similar exact sequences according to [27], and they are discussed in detail and used in the proof of Theorem 1.16 below.
3. The proofs of Theorems 1.4 and 1.7 and of Corollaries 1.8 and 1.10
3.1. Proof of Theorem 1.4
Observe that there is a natural $S$-linear morphism
@@
where $d=\deg C=e_{1}+e_{2}$ and $E$ denotes the Euler derivation, see [14, Theorem 5.1]. This morphism is injective. Indeed, if we suppose that $\theta(\delta)=0$ for some $\delta\in D_{0}(f_{1})_{k-e_{2}}$, then in particular $\theta(\delta)(f_{1})=0$. Since $E(f_{1})=e_{1}f_{1}\neq 0$ and $\delta(f_{1})=0$, we get $\delta(f_{2})=0$. It follows that
$0=\theta(\delta)=f_{2}\delta$
and hence $\delta=0$ since $f_{2}\neq 0$.
We denote by $d_{1},d_{2}$ (resp. by $d_{1}^{\prime},d_{2}^{\prime}$) the first two exponents of $C_{1}$ (resp. of $C$). By Proposition 1.3, we have
@@
This is a contradiction since the quotient $S$-module $D_{0}(f_{1})/Sr_{1}$ is torsion free, as it follows, for instance, from the exact sequence in [18, Theorem 5.1]. Indeed, this exact sequence may be rewritten as
$0\to S(-d_{1})\to D_{0}(f)\to B(C,r_{1})\to 0,$
where the first morphism is given by $h\mapsto hr_{1}$. Hence $D_{0}(f_{1})/Sr_{1}$ is isomorphic to an $S$-submodule II of the ideal $B(C,r_{1})$, in other words $I$ is an ideal in $S$, and hence has no torsion.
Remark 3.2.
Consider the Fermat curve $C_{1}:f_{1}=x^{d}+y^{d}+z^{d}=0$ and the line arrangement $C_{2}:f_{2}=x^{d}+y^{d}=0$. Then the curve $C_{1}$ has type d1d-1 since the exponents of $C_{1}$ are $d_{1}=d_{2}=d_{3}=d-1$. On the other hand, the union $C=C_{1}\cup C_{2}$ has type 0, since $C$ is a free curve, see [11, Corollary 1.6]. It follows that there is no interesting lower bound for the type of $C=C_{1}\cup C_{2}$ without providing any additional information on curves $C_{1}$ and $C_{2}$.
3.3. Proof of Theorem 1.7 and Corollaries 1.8 and 1.10
@@
and $r=e_{1}e_{2}$. Moreover, $D^{\prime}$ consists of $e_{2}$ points, and hence
@@
Example 3.4.
$1+2-2=1<t(C_{1})+\deg(C_{2})=2.$
Now we prove Corollary 1.8.
Proof.
The first claim follows from (1.3), since
@@
Moreover, (1.3) tells us that we can apply Theorem 1.7 to any line arrangement $C_{1}$. Starting with the integers $t$ and $d$ as in the second claim of Corollary 1.8, we choose $C_{1}$ to be a free line arrangement of $e_{1}=d-t\geq 3$ lines, for instance $x^{e_{1}}+y^{e_{1}}=0$. Then we add $t$ generic lines to this arrangement and we get by Theorem 1.7 a line arrangement $\mathcal{A}$} of degree $d$ and type $t$. This completes the proof of Corollary 1.8. ∎
Now we present our proof of Corollary 1.10.
Proof.
We first notice that our additional assumption on $\mathcal{C}$} implies
@@
for $e_{1}\geq 4$. To complete the proof of Corollary 1.10 it remains to add $t$ generic lines to the arrangement ${\mathcal{C}}_{1}$ and get in this way the required arrangement $\mathcal{C}$} by Theorem 1.7. ∎
@@
The curve $\mathcal{C}$} constructed above has a complicated singularity at the point $p=(0:0:1)$. Indeed, looking at $({\mathcal{C}},p)$ as the union of a $(e_{1}-2)$-ordinary point with the smooth germ at $p$ given by the conic $xz+y^{2}=0$ and using the known formula for the Milnor number of a union, see for instance [30, Theorem 6.5.1], one can show that the Milnor number $\mu({\mathcal{C}},p)$ is given by
@@
which was introduced in [8] and measures how far $({\mathcal{C}},p)$ is from being quasi-homogeneous, is given by
@@
4. The proof of Propositions 1.12, 1.13 and 1.14
4.1. Proof of Proposition 1.12
To see that the number $m$ of syzygies can be only $3$ or $4$, and to get the corresponding values for the $\varepsilon_{j}$’s, use the equality (1.7). In case (1), the value for $\tau(C)$ follows from [17, Proposition 2.1] using that $d_{1}+d_{2}=d+1$, where $d=\deg C$. In case (2), we use (1.6) to get the following resolution of the Milnor algebra $M(f)$:
@@
Here are some examples of curves $C$ having type $2$.
Example 4.2.
$f_{1}=(x^{2}+y^{2}-z^{2})(y-z)(x^{2}-z^{2})=0.$
$f_{1}=(-3x^{2}+xy+yz+xz)(-3y^{2}+xy+yz+zx)(-3z^{2}+xy+yz+zx)=0.$
This arrangement is known to be free with exponents $(2,3)$ and it has $3$ singularities of type $A_{5}$ and just one ordinary triple point $D_{4}$. If we add the conic $Q:f_{2}=x^{2}+y^{2}+z^{2}=0$, we get new conic arrangement $\mathcal{C}$ which has type $2A$ and exponents $(4,5,6)$. Note that the conic $Q$ is generic with respect to the arrangement $\mathcal{C}_{1}$, namely the intersection $\mathcal{C}_{1}\cap Q$ consists of $12$ points, which give rise to $12$ nodes for the arrangement $\mathcal{C}$.
Remark 4.3.
One very natural way to obtain line arrangements of type $2$ is the following. We start with a free line arrangement $\mathcal{A}$} and let ${\mathcal{A}}_{i,j}$ be the line arrangement obtained from $\mathcal{A}$} by deleting two lines, namely $L_{i}$ and $L_{j}$. Then [4, Theorem 1.8] gives sufficient condition such that $t({\mathcal{A}}_{i,j})=2$. More precisely, with the notation from this result, we have the following possibilities. The arrangements in [4, Theorem 1.8] (1) have type $2B$, and the same holds for the arrangements in case (2), when $c_{1}<c_{2}$. On the other hand, the arrangements in [4, Theorem 1.8] (2) have type $2A$ if $c_{1}=c_{2}$ and $t({\mathcal{A}}_{i,j})\neq 1$. Moreover, in [4, Example 4.5], the author produces a free line arrangement $\mathcal{A}$} of $11$ lines, such that the arrangement ${\mathcal{A}}_{3,4}$ has type $2A$ and the arrangements ${\mathcal{A}}_{2,11}$ and ${\mathcal{A}}_{3,5}$ have type $2B$.
4.4. Proof of Proposition 1.13 and 1.14
We start with Proposition 1.13. Here we use Theorem 2.1 and Proposition 1.12. Note that the condition $d_{1}<d_{2}-2$ is exactly the condition $d_{1}<(d-1)/2$ which occurs in Theorem 2.1 (1). Hence, to prove the claim $(1)$ in Proposition 1.13, we use the formula for $\nu(C)$ given in Theorem 2.1 (1) and the formula for $\tau(C)$ given in Proposition 1.12, $(1)$. To prove the claims (2), (3) and (4), we use the formula for $\nu(C)$ given in Theorem 2.1 (2) and the formula for $\tau(C)$ given in Proposition 1.12 $(1)$. This completes the proof of Proposition 1.13.
@@
5. Proof of Theorem 1.15, Theorem 1.16, Proposition 1.18 and Theorem 1.20
5.1. Proof of Theorem 1.15
We use the exact sequence (2.6) above with $C_{1}={\mathcal{A}}^{\prime}:f_{1}=0$ and $C_{2}=L:f_{2}=0$. Hence $e_{1}=m_{1}+m_{2}=r$, $e_{2}=1$, $g_{2}=\operatorname{genus}(C_{2})=0$. Moreover, $D^{\prime}$ is just a point, and hence
@@
Let $(d_{1},d_{2},\ldots,d_{m})$ be the exponents of the curve $C={\mathcal{A}}$.
First, consider the case $k<m_{1}+m_{2}-1=e_{1}-1$. Then $H^{0}(L,{\mathcal{O}}_{L}(k+1-m_{1}-m_{2}))=0$ and we get an isomorphism
@@
5.2. Proof of Theorem 1.16
This rather long proof is divided into two steps to help the reader understand it more easily.
Step 1. We use again the exact sequence (2.6) above with $C_{1}={\mathcal{C}}^{\prime}:f_{1}=0$ and $C_{2}=Q:f_{2}=0$. Hence we have $e_{1}=m_{1}+m_{2}+1$, $r=2e_{1}$, $e_{2}=2$, $g_{2}=\operatorname{genus}(C_{2})=0$. Moreover, $D^{\prime}$ consists of $2$ points and hence
@@
Let $(d_{1},d_{2},\ldots,d_{m})$ be the exponents of the curve $C={\mathcal{C}}$.
First, consider the case $k<e_{1}$. Then $H^{0}(L,{\mathcal{O}}_{L}(2k-2e_{1}))=0$ and we get an isomorphism
@@
Step 2. If the conic-line arrangement $\mathcal{C}$} has type $2A$, then the exponents of $\mathcal{C}$} are $(d_{1},d_{2},d_{3})=(m_{1}+2,m_{2}+2,m_{1}+m_{2}+1)$, and our result is proved. To show that this is the case, we have to show that there is no need for an additional generator for $D_{0}(f)$ beyond $\alpha$}.
A triple $({\mathcal{C}}^{\prime},{\mathcal{C}},{\mathcal{C}}^{\prime\prime})$ of CL (short for conic-line) arrangements, where ${\mathcal{C}}^{\prime}={\mathcal{C}}\setminus\{Q\}$ with $Q\in{\mathcal{C}}$ a smooth conic and $\mathcal{C}$$^{\prime\prime}$ the 0-dimensional arrangement on $Q$ given by the reduced intersection ${\mathcal{C}}^{\prime}\cap Q$, is called quasi-homogeneous if all the singular points of $\mathcal{C}$$^{\prime}$ and of $\mathcal{C}$} situated on $Q$ are quasi-homogeneous.
Theorem 5.3.
If a triple $({\mathcal{C}}^{\prime},{\mathcal{C}},{\mathcal{C}}^{\prime\prime})$ of CL is quasi-homogeneous, then the associated exact sequence (2.5) with $_{1}={\mathcal{C}}^{\prime}, C=𝒞$$\mathcal{C}$} and $C_{2}=Q$ is exact.
Proof.
This result is a stronger version of [27, Proposition 3.7]. The authors in [27] require in their result that all the singular points of $\mathcal{C}$$^{\prime}$ and of $\mathcal{C}$} are quasi-homogeneous, but this stronger assumption is not necessary. Indeed, as noted in [14] and in subsection 2.2 above, for a similar condition in [28], this assumption is used in [27] only to compute the difference
@@
[…]
see for instance the proof of [27, Proposition 2.8 and Proposition 3.7]. The difference $\tau({\mathcal{C}})-\tau({\mathcal{C}}^{\prime})$ can be expressed in the following way
$\sum_{p\in{\mathcal{C}}^{\prime\prime}}(\tau({\mathcal{C}},p)-\tau({\mathcal{C}}^{\prime},p)),$
since for a point $q\in{\mathcal{C}}\setminus{\mathcal{C}}^{\prime\prime}$ one has $\tau({\mathcal{C}},q)-\tau({\mathcal{C}}^{\prime},q)=0$, as the two singularities $({\mathcal{C}},q)$ and $\tau({\mathcal{C}}^{\prime},q)$ coincide for $q\notin Q$, $Q$ is smooth and we set by convention $\tau({\mathcal{C}}^{\prime},p)=0$ when the germ $({\mathcal{C}}^{\prime},p)$ is the empty germ. Lemma 2.7 in [27] is applied to the singularities $p\in{\mathcal{C}}^{\prime\prime}$, since for quasi-homogeneous singularities the Milnor number coincide with the Tjurina number. That is why it is enough to require that the singularities of $\mathcal{C}$} and $\mathcal{C}$$^{\prime}$ situated on $Q$ are quasi-homogeneous. ∎
@@
Using this isomorphism, we can regard the 0-dimensional arrangement $\mathcal{C}$$^{\prime\prime}$ as a 0-dimensional arrangement in $\mathbb{P}^{1}$ and hence it is defined by an equation $f^{\prime\prime}=0$, where $f^{\prime\prime}\in R=\mathbb{C}[s,t]$ is some homogeneous polynomial. Hence the graded $R$-module $D(f^{\prime\prime})$ is well-defined. Moreover, the map $v$ above induces a ring morphism
@@
and the $S$-module $D(f^{\prime\prime})_{\psi}$ is just the module obtained from the $R$-module $D(f^{\prime\prime})$ by restriction of scalars using $\psi$. Next, we consider the morphisms in the exact sequence (5.3). The first morphism $D(f^{\prime})(-2)\to D(f)$ is given by the multiplication by $q$, the equation of $Q$, and hence it preserves the homogeneous components, that is
@@
After these preliminaries, we can continue our proof in Step 2. In our situation, $\mathcal{C}$$^{\prime\prime}$ consists of $2e_{1}$ points, hence $f^{\prime\prime}$ is a homogeneous polynomial of degree $2e_{1}$ without multiple factors. The derivation ${\alpha}\in D_{0}(f)_{e_{1}}$ that was constructed at the end of Step 1 can be regarded as sitting in $\overline{D}(f)_{e_{1}}$ by using the identification from Remark 2.3. Then any representative ${\alpha}_{0}\in D(f)_{e_{1}}$ of $\alpha$} is clearly not in the image of the morphism $D(f^{\prime})(-2)\to D(f)$ and hence $\rho({\alpha}_{0})\neq 0$. If we assume ${\alpha}_{0}=(a_{1},a_{2},a_{3})$, then
@@
would be a new representative of $\alpha$} and clearly $\rho({\alpha}_{1})=0$, a contradiction.
If $\mathcal{C}$} has type $2B$, it means that there is an additional generator of degree $d_{4}>d_{3}=e_{1}$. Using (1.4), we get $d_{4}\leq(e_{1}+2)-1$, and hence $d_{4}=e_{1}+1$. On the other hand, the exact sequence (5.2) which holds due to Theorem 5.3, where $e_{1}$ is replaced by $e_{1}+1$, shows that $\overline{D}_{0}(f)_{e_{1}+1}$ modulo the image of $\overline{D}_{0}(f_{1})_{e_{1}-1}$ is 2 dimensional. It follows that we need a new generator of degree $e_{1}+1$ if and only the vector space $V$ spanned by $x{\alpha}_{0},y{\alpha}_{0}$ and $z{\alpha}_{0}$ in $\overline{D}_{0}(f)_{e_{1}+1}$ modulo the image of $\overline{D}(f_{1})_{e_{1}-1}$ is 1-dimensional. Note that $\rho(E)=E^{\prime\prime}$, the Euler derivation corresponding to the ring $R$, and hence there is an induced morphism
@@
Example 5.4.
An illustration of Theorem 1.16 occurs in Example 4.2 (iii), where $(m_{1},m_{2})=(2,3)$. Note that in Example 4.2 (ii) the condition that the intersection $Q\cap{\mathcal{A}}^{\prime}$ consists of $2\deg{\mathcal{A}}^{\prime}=2(m_{1}+m_{2}+1)=10$ points is not fulfilled, since $\tau(\mathcal{CL})=24$ by Proposition 1.12, and hence
$\tau(\mathcal{CL}_{1})+10=22\neq\tau(\mathcal{CL}).$
This explains why the type $2B$ may occur here.
5.5. Proof of Proposition 1.18
@@
5.6. Proof of Theorem 1.20
This proof is similar to our proof of Theorem 1.16, so we divide it into two steps.
@@
Let $(d_{1},d_{2},\ldots,d_{m})$ be the exponents of the curve $C={\mathcal{A}}$.